# Test-Time Canonicalization by Foundation Models for Robust Perception

### v1

@@ Abstract
Real-world visual perception requires invariance to diverse transformations, yet current methods rely heavily on specialized architectures or training on predefined augmentations, limiting generalization. We propose FoCal, a test-time, data-driven framework that achieves robust perception by leveraging internet-scale visual priors from foundation models. By generating and optimizing candidate transformations toward visually typical, “canonical” views, FoCal enhances robustness without retraining or architectural changes. Experiments demonstrate improved robustness of CLIP and SAM across challenging transformations, including 2D/3D rotations, illumination shifts (contrast and color), and day-night variations. We also highlight potential applications in active vision. Our approach challenges the assumption that transform-specific training is necessary, instead offering a scalable path to invariance. Our code is available at: https://github.com/sutkarsh/focal.
@@
Figure 1: Introducing a test-time approach to invariance for complex and diverse transforms: We propose FoCal, a general test-time framework for achieving approximate invariance to complex transformations at scale. Given variations in viewpoint, illumination, environment, and rotation, FoCal leverages foundation model priors to select a visually-typical version of the input. As depicted in the bottom, out-of-distribution inputs can lead to incorrect predictions on CLIP (“bench”) and incorrect segmentations on SAM (not getting the whole vase) that are corrected when running CLIP and SAM on the FoCal canonicalized versions. Thus, FoCal offers a scalable, data-driven method for robust perception across diverse transforms.
@@ 1 Introduction
A robot navigating a cluttered home or a self-driving car cruising down a street must perceive objects consistently across many physical transformations – fluctuating viewpoints, orientations, lighting conditions, and more. This capability, known as invariant visual perception, separates fragile algorithms from robust real-world systems.
However, despite being trained on hundreds of millions of images, even leading foundation models such as OpenAI’s CLIP (Radford et al., 2021) are brittle against such transformations (Figure 1) – CLIP misclassifies images with unexpected viewpoint shifts, and SAM (Kirillov et al., 2023) fails to produce the correct segmentation on unexpectedly rotated images. One hypothesis for this brittleness is that internet-scale training data overrepresents certain poses and lighting most commonly encountered by humans in day-to-day life, leaving models unprepared for atypical (or out-of-distribution) inputs often seen in broader embodied settings (Madan et al., 2024). This phenomenon is sometimes referred to as the photographer’s bias.
Data augmentation (DA) and equivariant networks have thus emerged as valuable tools to improve robustness to out-of-distribution transformations. DA trains models on carefully designed augmentations (rotations, crops, etc.) (Bouchacourt et al., 2021), while equivariant networks (Cohen & Welling, 2016) bake mathematical symmetries (e.g., 2D rotations) into architectures. These approaches succeed when transformations are known a priori and well-sampled during training. However, they face inherent limitations in open-world scenarios: DA struggles with rare classes (Zhou et al., 2022) and distribution shifts unseen during training, while equivariant architectures cannot handle complex real-world transforms like 3D viewpoint shifts. Importantly, both methods require designers to anticipate all possible transformations during training, a difficult requirement for systems operating in diverse environments.
We instead propose Foundation-model guided Canonicalization (FoCal), a training-free framework that reasons about transforms through test-time optimization (Figure 1). Our key insight is that the visual priors learned by foundation models can guide us toward the most visually-typical view of any input, much like how humans mentally transform objects to familiar orientations before recognition (Shepard & Metzler, 1971). FoCal thus follows a “vary and rank” scheme (Figure 2): it produces candidate transforms (e.g., novel views using a generative model like Xiang et al. (2024)) and selects the most visually typical instance by minimizing an energy function based on CLIP and Stable Diffusion. These “canonicalized” views are then passed on to downstream models.
@@
This optimization-driven approach enables robust visual perception across a wide range of challenging conditions, including changes in viewpoint, lighting, and environment, without requiring additional training or architectural changes. Unlike prior canonicalization approaches that require transform-specific training (Mondal et al., 2023), FoCal works for diverse transforms and datasets. Our method is similar to the recent test-time compute scaling work that uses a learned ranker to pick the best among many outputs (Snell et al., 2024). FoCal reframes invariance as a search over the implicit priors encoded in foundation models, thus offering a scalable, data-driven method for robust perception across diverse transforms.
@@
We evaluate FoCal on 3D-viewpoint, illumination, and day-night changes, as well as 2D rotations. We find that FoCal successfully canonicalizes these transforms, improving out-of-distribution performance of downstream foundation models such as CLIP (Radford et al., 2021) on ImageNet (Deng et al., 2009) scale datasets. To our knowledge, this is a significant improvement in dataset size and transformation complexity compared to previous canonicalization work. Remarkably, FoCal consistently matches or outperforms task-specific canonicalizers like PRLC (Mondal et al., 2023) even in their trained settings, despite being a completely test-time approach.
@@
Contributions: (1) FoCal: A test-time, data-driven framework using the visual priors of foundation models for invariance; (2) An approximate invariance method that scales to complex transforms such as 3D viewpoint shifts, lighting changes, and environmental variations; (3) Evaluations on modern models such as CLIP, OV-Seg, and SAM on datasets like ImageNet, COCO, Objaverse-LVIS, and CO3D.
@@
Our results challenge the prevailing assumption that invariance requires specific training or architectures, offering a path toward robust perception for embodied agents. By bridging insights from human mental rotation and modern test-time compute scaling (Zaremba et al., 2025), FoCal shows that test-time optimization can be used for general, scalable, and physically robust perception.
@@ 2.1 Problem setting
[…]
@@
where $E:\mathcal{X}\to\mathbb{R}$ is a real-valued function referred to as the “energy function”. Strikingly, this result does not require $E$ or $\bm{f}$ to be equivariant or invariant themselves. Instead, the canonicalizer $\bm{h}$ emerges from the minimization process over the energy function $E$.
@@ 2.3 Energy Functions from Foundation Models
where $\mathbf{x_{t}}=\sqrt{\bar{\alpha}_{t}}\bm{x}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ is the noisy input, $\bm{\epsilon_{\theta}}$ is the pre-trained diffusion model, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ with denoising schedule $\beta_{t}$. We use SD-2-base (Rombach et al., 2021), finding that $5-10$ steps are usually sufficient.
@@ 2.4 Foundation Model Guided Canonicalization
‘Vary’ and ‘Rank’: Our method consists of two steps: Vary, where we produce variations of an input image, and Rank, where we rank the variations (Figure 2). More formally, we optimize the FoCal energy function over transformations of the input image $\bm{x}$:
@@ 3 Experiments
We evaluate FoCal as a method for approximate invariance to viewpoint transforms on Objaverse-LVIS (Deitke et al., 2023) and CO3D (Reizenstein et al., 2021). Objaverse-LVIS dataset contains $46$K 3D assets and CO3D contains  19K real multiview video sequences with object segmentations.
We begin by filtering datasets to ensure high-quality objects and viewpoints. For Objaverse-LVIS, we noticed cases of misleading and overlapping labels and thus filtered out such objects. For CO3D, we use sequences with sufficient variation in viewpoint quality. Given the filtered videos, we then sample one of the poor viewpoints (i.e., ground truth probability of $<t$ for some threshold $t$) and filter out images with blurriness or segmentation errors (Appendix B.1).
We use TRELLIS (Xiang et al., 2024) to produce viewpoint variations for each input. Specifically, we render views on the sphere at $30^{\circ}$ intervals for the “vary” stage (Appendix B.1). We rank these renders using FoCal energy and pick the view with the minimum energy. Since TRELLIS generates background-removed images, we use OV-Seg (Liang et al., 2023), a version of CLIP (Radford et al., 2021) fine-tuned on background-removed images, as our downstream classifier.
@@ 3.2 Lighting (color and contrast)
Figure 8: FoCal beats PRLC on their image classification evaluations: We evaluate FoCal on 2D rotations following the setting of Mondal et al. (2023). (a) Left plot shows classification accuracy gains relative to rotation augmentation baseline across multiple datasets (CIFAR10, CIFAR100, STL10) using both ResNet-50 and ViT architectures. Each paired line represents a different evaluation setting, with FoCal (Ours) consistently matching or outperforming PRLC in each setting. (b) We evaluate each of PRLC’s trained canonicalizers on an unseen dataset (ImageNet-1K) and pick the best canonicalizer for each architecture. We find that FoCal still beats them as PRLC’s canonicalizers struggle to generalize outside their training setting. This showcases FoCal’s generalizability across datasets. In short, FoCal outperforms trained canonicalizers of Mondal et al. (2023) on seen as well as unseen datasets.
@@
We extend this setting further to CLIP and ImageNet (Deng et al., 2009), finding that FoCal generalizes much better to this setting (by up to $+11\%$) than PRLC, which struggles to generalize beyond its training scope (Figures 8 and 5).
@@
Figure 10: Applying FoCal to active vision: We optimize a 6-DoF camera pose (x, y, z, yaw, pitch, roll) in a 3D virtual environment. We use Bayesian optimization to minimize FoCal energy, and plot the results above along with randomly chosen poses for comparison. We find that the camera naturally focuses on salient objects and tends to maintain upright angles, reflecting behaviors observed in simpler settings (green underline). Although some failure cases remain (red underline), these results illustrate FoCal ’s ability to generalize to significantly more complex transformations and viewing conditions in a zero-shot manner.
Segmentation: We also evaluate PRLC’s segmentation setting with SAM (Kirillov et al., 2023) using the pre-trained SAM canonicalizer supplied by the authors. Here, we find that our method matches their pre-trained canonicalizer in mAP ($65.9$) while achieving $+2.1\%$ better pose accuracy.
@@ 3.5 Active vision
We pick $8$ scenes (Figure 10) and use the same translation ($[-3,3]$) and rotation range (full rotations) for each scene. We use ImageNet classes to compute CLIP energy and minimize it using $150$ iterations of BO. We also found that using $450$ initial random points can help exploration but is unnecessary for most scenes. We hope to study the optimization schedules more systematically in the future.
@@ 4 Related Work
These developments have inspired progress in deep learning architectures (Jaderberg et al., 2015; Boominathan et al., 2016). More recently, Kaba et al. (2022) propose a learned canonicalization (LC) approach via minimizing a learned energy function. At test time, it minimizes this function to canonicalize the inputs before passing them through the downstream model, enabling better generalization. PRLC (Mondal et al., 2023) adds a regularization prior to align the training set distributions with the predicted canonical for better accuracy.
@@
3D robustness and pose estimation: Existing approaches for 3D robustness pool features across multiple views (Fan et al., 2024; Su et al., 2015; Yang & Wang, 2019; Wei et al., 2020, 2022; Hamdi et al., 2021; Kanezaki et al., 2018), whereas our technique only requires one view at test time. Chen et al. (2020) learns category-level pose estimation using analysis-by-synthesis. This approach is related to our approach; however, it is category-specific, whereas our model is category-agnostic. ImageNet3D (Ma et al., 2024) annotates a large dataset of 3D objects with poses and trains NNs for open-set pose estimation in a supervised manner. In contrast, our method is unsupervised.
@@ 5 Limitations and Future Work
Speeding up FoCal. Evaluating the energy function for many candidates is computationally expensive (Section A.10). Specifically, the complexity is $N\times\left(C_{\text{transform}}+C_{\text{energy}}+C_{\text{inference}}\right)$, where $N$ is the number of transforms. This limitation is similar to recent LLMs that use test-time search for better reasoning and robustness. Future work could explore a System-1/2 scheme where canonicalization is only used when necessary, enabling users to gain the robustness benefits of FoCal without incurring as much of a cost in runtime efficiency. As a preliminary experiment, in 2D, we were able to detect upright vs. non-upright images with 95% accuracy by comparing the input against $+90^{\circ}$ and $-90^{\circ}$ and thresholding the energy difference.
@@ 6 Conclusion
    • Cross-Dataset Analysis: Shows $<3\%$ variance in pose error vs PRLC’s 12–18% drops (Figure 14)
@@ A.4 Segmentation Results
Table 4: FoCal matches PRLC on segmentation: We first report FoCal’s performance on mAP on COCO, PRLC without any training (left). We then show that the $C4$ pose accuracy is higher than PRLC by $2.1\%$. This shows FoCal’s ability to generalize to segmentation without supervision. All numbers are reproduced using Mondal et al. (2023)’s pre-trained checkpoint.
@@ A.6 Cross-Dataset and Model Analysis
FoCal ’s unified framework enables superior cross-domain transfer compared to PRLC’s (Mondal et al., 2023) specialized canonicalizers. When transferring across datasets, PRLC suffers 12-18% drops in pose accuracy (LABEL:fig:data_transfer_r50_acc), while our method maintains $<3\%$ variance in pose error. Transferring to CLIP (Radford et al., 2021) yields particularly strong results, with +15% relative accuracy gains over PRLC (Figure 15). These results underscore a key advantage: by avoiding dataset-specific training, FoCal develops general canonicalization strategies that transfer seamlessly across both architectures (ResNet50/ViT/CLIP) and data distributions (CIFAR/STL10/ImageNet).

### v2

@@ Abstract
Perception in the real world requires robustness to diverse viewing conditions. Existing approaches often rely on specialized architectures or training with predefined data augmentations, limiting adaptability. Taking inspiration from mental rotation in human vision, we propose FoCal, a test-time robustness framework that transforms the input into the most typical view. At inference-time, FoCal explores a set of transformed images and chooses the one with the highest likelihood under foundation model priors. This test-time optimization boosts robustness while requiring no retraining or architectural changes. Applied to models like CLIP and SAM, it significantly boosts robustness across a wide range of transformations, including 2D and 3D rotations, contrast and lighting shifts, and day-night changes. We also explore potential applications in active vision. By reframing invariance as a test-time optimization problem, FoCal offers a general and scalable approach to robustness. Our code is available at: https://github.com/sutkarsh/focal.
@@
Figure 1: Introducing FoCal: a test-time framework for approximate invariance to complex transformations at scale. (top) Given variations in viewpoint, illumination, environment, and rotation, FoCal use foundation model priors to select a visually-typical version of the input. (bottom) out-of-distribution inputs can lead to incorrect predictions on CLIP (“bench”) and incorrect segmentations on box-prompted SAM (missing body parts in the segmentation mask). Using FoCal canonicalized versions fixes these errors. Thus, FoCal offers a scalable, data-driven method for robust perception across diverse transforms.
@@ 1 Introduction
A robot navigating a cluttered home or a self-driving car on a street must reliably recognize objects despite changes in viewpoint, orientation, and lighting. Such robustness is essential for embodied perception in dynamic environments.
Human vision has evolved to excel at this task: We readily recognize familiar objects from diverse viewpoints while mentally rotating unfamiliar objects to typical views to facilitate recognition (Shepard & Metzler, 1971). Through natural visual exposure, we develop invariances that adapt to each object and context, generalizing to novel ones while preserving necessary distinctions. That is, human perception develops invariances shaped by ecological demands.
In contrast, even leading foundation models exhibit brittleness (Figure 1): CLIP (Radford et al., 2021) misclassifies images from uncommon viewpoints, and SAM (Kirillov et al., 2023) incorrectly segments objects presented sideways. This vulnerability stems partly from the overrepresentation of upright poses and ideal lighting in internet training data – a phenomenon known as photographer’s bias. Consequently, these models are unprepared for diverse visual inputs seen in embodied settings (Madan et al., 2024).
Data augmentation (DA) and equivariant networks are two common approaches to improve neural network robustness to specific transformations. DA exposes models to carefully selected transformations during training (e.g., rotations, crops), but struggles with rare classes (Zhou et al., 2022) and hurts some classes through over-regularization (Xiao et al., 2021). Equivariant networks (Cohen & Welling, 2016) build mathematical symmetries (e.g., 2D rotations) directly into the architecture, but this approach does not extend to complex real-world transformations like 3D viewpoint changes.
Critically, both approaches hard-wire invariances into the model during training through either curated training data or the model architecture. This strategy is inherently rigid: it assumes prior knowledge of relevant transformations and often fails under distribution shifts such as novel viewpoints or lighting conditions beyond those seen during training.
We propose a fundamentally different approach: Instead of hard-wiring invariance during training, our Foundation-model guided Canonicalization (FoCal) achieves invariance by reasoning over transformations at inference time (Figure 1). Like human vision, which mentally rotates unfamiliar objects to canonical views for recognition (Shepard & Metzler, 1971), FoCal leverages foundation models’ visual priors to transform inputs toward visually typical views.
FoCal uses a “vary and rank” strategy (Figure 2): It generates candidate transforms (e.g., via generative models) and selects the most visually typical instance by minimizing an energy function based on Stable Diffusion (SD) and CLIP. The optimized instance is then passed to downstream models for inference. This perspective connects to recent test-time scaling work where learned rankers select among outputs (Snell et al., 2024). However, FoCal distinguishes itself by grounding ranking in data-driven visual typicality.
@@
Our approach enables robust perception across diverse conditions, including changes in viewpoint, lighting, and environment, without any additional training or architectural changes. Invariance in our method does not arise from hard-wired constraints, but is instead an emergent property of our test-time optimization. Unlike prior canonicalization work that requires transform-specific training (Mondal et al., 2023), FoCal works for a wide range of datasets and transforms. By reframing invariance as test-time optimization, FoCal offers a data-driven method for robustness at scale.
@@
We evaluate FoCal across a range of challenging transformations, including 3D-viewpoint, illumination, day-night changes, and 2D rotations. We find that FoCal improves out-of-distribution performance of foundation models such as CLIP (Radford et al., 2021) on ImageNet (Deng et al., 2009) scale datasets. To our knowledge, this is a significant improvement in dataset size and transformation complexity compared to previous canonicalization work. Remarkably, FoCal consistently matches or outperforms task-specific canonicalizers like PRLC (Mondal et al., 2023) even in their trained settings, despite operating entirely at test time.
@@
Contributions. (1) We introduce FoCal, a test-time, data-driven framework that leverages the visual priors of foundation models to achieve invariance. (2) We propose an approximate invariance method that scales to complex transformations, including 3D viewpoint shifts, lighting changes, and environmental variations. (3) We demonstrate the effectiveness of FoCal through evaluations on modern models such as CLIP, OV-Seg, and SAM, across diverse datasets including ImageNet, COCO, Objaverse-LVIS, and CO3D.
@@
Our results challenge the prevailing assumption that achieving invariance requires specialized training or architectural design, offering a path toward robust perception for embodied agents. By bridging insights from human mental rotation and modern test-time compute scaling (Zaremba et al., 2025), FoCal demonstrates that test-time optimization can serve as a scalable approach to robust visual perception under real-world variation.
@@ 2.1 Problem setting
$\text{{Equivariance:}}\quad\bm{f}(t(\bm{x}))=t^{\prime}(\bm{f}(\bm{x}))\ \ \forall t\in\mathcal{T}$
@@
where $E:\mathcal{X}\to\mathbb{R}$ is a real-valued function referred to as the “energy function”. Strikingly, this result does not require $E$ or $\bm{f}$ to be equivariant or invariant. Instead, the canonicalizer $\bm{h}$ emerges from the minimization process.
@@ 2.3 Energy Functions from Foundation Models
where $\mathbf{x_{t}}=\sqrt{\bar{\alpha}_{t}}\bm{x}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ is the noisy input, $\bm{\epsilon_{\theta}}$ is the pre-trained diffusion model, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ with denoising schedule $\beta_{t}$. We use SD-2-base (Rombach et al., 2021), finding that 5105-10 steps are usually sufficient.
@@ 2.4 Foundation Model Guided Canonicalization
‘Vary’ and ‘Rank’: Our method consists of two steps: Vary, where we produce variations of an input image, and Rank, where we rank the variations (Figure 2). We optimize the FoCal energy function over transforms of the input $\bm{x}$:
@@ 3 Experiments
We evaluate FoCal as a method for approximate invariance to viewpoint transforms on Objaverse-LVIS (Deitke et al., 2023) and CO3D (Reizenstein et al., 2021). Objaverse-LVIS dataset contains 4646K 3D assets and CO3D contains  19K real multiview video sequences with object segmentations.
We begin by filtering datasets to ensure high-quality objects and viewpoints. For Objaverse-LVIS, we noticed cases of misleading and overlapping labels and thus filtered out such objects. For CO3D, we use sequences with sufficient variation in viewpoint quality. Given the filtered videos, we then sample one of the poor viewpoints (i.e., ground truth probability of <t<t for some threshold $t$) and filter out images with blurriness or segmentation errors (Appendix B.1).
We use TRELLIS (Xiang et al., 2024) to produce viewpoint variations for each input. Specifically, we render views on the sphere at $30^{\circ}$ intervals for the “vary” stage (Appendix B.1). We rank these renders using FoCal energy and pick the view with the minimum energy. Since TRELLIS generates background-removed images, we use OV-Seg (Liang et al., 2023), a version of CLIP fine-tuned on background-removed images, as our classifier.
@@ 3.2 Lighting (color and contrast)
Figure 8: FoCal beats PRLC on their image classification evaluations: We evaluate FoCal on 2D rotations following the setting of Mondal et al. (2023). (a) Left plot shows classification accuracy gains relative to rotation augmentation baseline across multiple datasets (CIFAR10, CIFAR100, STL10) using both ResNet-50 and ViT architectures. Each paired line represents a different evaluation setting, with FoCal (Ours) consistently matching or outperforming PRLC in each setting. (b) We evaluate each of PRLC’s trained canonicalizers on an unseen dataset (ImageNet-1K) and pick the best canonicalizer for each architecture. We find that FoCal still beats them as PRLC’s canonicalizers struggle to generalize outside their training setting. In short, FoCal outperforms trained canonicalizers of Mondal et al. (2023) on seen as well as unseen datasets, highlighting FoCal’s generalizability.
@@
We extend this setting further to CLIP and ImageNet (Deng et al., 2009), finding that FoCal generalizes much better to this setting (by up to $11\%$) than PRLC, which struggles to generalize beyond its training scope (Figures 8 and 5).
@@
Figure 10: Applying FoCal to active vision: We optimize a 6-DoF camera pose (x, y, z, yaw, pitch, roll) in a 3D virtual environment. We use Bayesian optimization to minimize FoCal energy, and plot the results above along with randomly chosen poses for comparison. The first row shows the initialization, second row shows a random pose, and the third row shows the result from FoCal. We find that the camera naturally focuses on salient objects and tends to maintain upright angles, reflecting behaviors observed in simpler settings (green underline). Although some failure cases remain (red underline), these results illustrate FoCal ’s ability to generalize to significantly more complex transformations and viewing conditions in a zero-shot manner.
Segmentation: We also evaluate PRLC’s segmentation setting with SAM (Kirillov et al., 2023) using the pre-trained SAM canonicalizer supplied by the authors. Here, we find that our method matches their pre-trained canonicalizer in mAP ($65.9$) while achieving $2.1\%$ better pose accuracy.
@@ 3.5 Active vision
We pick $8$ scenes (Figure 10) and use the same translation ([3,3][-3,3]) and rotation range (full rotations) for each scene. We use ImageNet classes to compute CLIP energy and minimize it using $150$ iterations of BO. We also found that using $450$ initial random points can help exploration but is unnecessary for most scenes. We hope to study the optimization schedules more systematically in the future.
@@ 4 Related Work
These developments have inspired progress in deep learning architectures (Jaderberg et al., 2015; Boominathan et al., 2016). More recently, Kaba et al. (2022) propose a learned canonicalization (LC) approach via minimizing a learned energy function. At test time, it minimizes this function to canonicalize the inputs. PRLC (Mondal et al., 2023) adds a regularization prior to align the canonical with the train set distribution for better accuracy.
@@
3D robustness and pose estimation: Existing approaches for 3D robustness pool features across multiple views (Fan et al., 2024; Su et al., 2015; Yang & Wang, 2019; Wei et al., 2020, 2022; Hamdi et al., 2021; Kanezaki et al., 2018), whereas our technique only requires one view at test time. Chen et al. (2020) learns category-level pose estimation using analysis-by-synthesis. This approach is related to our approach; however, it is category-specific, whereas our model is category-agnostic. ImageNet3D (Ma et al., 2024) annotates a large dataset of 3D objects with poses for open-set pose estimation, whereas our method is unsupervised.
@@ 5 Limitations and Future Work
Speeding up FoCal. Evaluating the energy function for many candidates is computationally expensive (Section A.10). Specifically, the complexity is $N\times\left(C_{\text{transform}}+C_{\text{energy}}+C_{\text{inference}}\right)$, where $N$ is the number of transforms. This limitation is similar to recent LLMs that use test-time search for better reasoning and robustness. Future work could explore a System-1/2 scheme where canonicalization is only used when necessary, enabling users to gain the robustness benefits of FoCal without incurring as much of a cost in runtime efficiency. As a preliminary experiment, in 2D, we were able to detect upright vs. non-upright images with 95% accuracy by comparing the input against $90^{\circ}$ and $90^{\circ}$ and thresholding the energy difference.
@@ 6 Conclusion
    • Cross-Dataset Analysis: Shows $3\%$ variance in pose error vs PRLC’s 12–18% drops (Figure 14)
@@ A.4 Segmentation Results
Table 4: FoCal matches PRLC on segmentation: We first report FoCal’s performance on mAP on $CO$, PRLC without any training (left). We then show that the $C4$ pose accuracy is higher than PRLC by $2.1\%$. This shows FoCal’s ability to generalize to segmentation without supervision. All numbers are reproduced using Mondal et al. (2023)’s pre-trained checkpoint.
@@ A.6 Cross-Dataset and Model Analysis
FoCal ’s unified framework enables superior cross-domain transfer compared to PRLC’s (Mondal et al., 2023) specialized canonicalizers. When transferring across datasets, PRLC suffers 12-18% drops in pose accuracy (LABEL:fig:data_transfer_r50_acc), while our method maintains $3\%$ variance in pose error. Transferring to CLIP (Radford et al., 2021) yields particularly strong results, with +15% relative accuracy gains over PRLC (Figure 15). These results underscore a key advantage: by avoiding dataset-specific training, FoCal develops general canonicalization strategies that transfer seamlessly across both architectures (ResNet50/ViT/CLIP) and data distributions (CIFAR/STL10/ImageNet).